OpenStudy (anonymous):

he profits in a business are to be shared by three partners in the ratio of 3 to 2 to 5. The profits for the year was 176,500.determine the number of dollars each partner is to receive. @mitu12

OpenStudy (mitu12):

If the partners share the profits in the ratio 2:3:5, then for every $2 + $3 + $5 = $10 of profit, the first partner receives $2, the second partner receives $3, and the third partner receives $5. Thus the first partner gets $2 / $10 = 1 / 5, or .2, of the total. This means he gets $176500 x .2 = $35300. The second partner gets $3 / $10 = 3 /10 = .3 of the total, so he gets $176500 x .3 = $52950. The third partner gets $5 / $10 = 1 / 2, or .5 of the total. Then he gets $176500 x .5 = $88250. To check, notice that $35300 + $52950 + $88250 = $176500. When you add the three shares together, you get the total profit for the year!

OpenStudy (anonymous):

That's all ?

OpenStudy (mitu12):

yes

OpenStudy (mitu12):

do you understand it?

OpenStudy (anonymous):

Yes, is it okay if you would help me with my other question

OpenStudy (mitu12):

sure

OpenStudy (anonymous):

Thanks